what is the length and the width of a recatangle with an area of 38,646 sq ft?

Answer by Alan3354(69443)

what is the length and the width of a recatangle with an area of 38,646 sq ft?
--------------
Area = L*W
L*W = 38646
L = 38646/W
-----
Pick any value for W>0, find L